OpenStudy (anonymous):

the solutions to the quadratic equation x^2+5x+4=0 are A=0 and 1 B=-1and4 C.-1and-1 D.3 and 4 E.there are no real roots

OpenStudy (anonymous):

@FriedRice

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

well you can just simplify the equation

OpenStudy (anonymous):

or you can put the equation into the quadratic formula which is \[-b \pm \frac{ \sqrt b^2 - 4ac }{ 2a }\]

OpenStudy (anonymous):

so what would the answer be??

OpenStudy (anonymous):

a = 1 b = 5 c = 4

OpenStudy (anonymous):

would the answer be B

OpenStudy (anonymous):

ok is the answers correct?

OpenStudy (anonymous):

i dont see the right answer

OpenStudy (anonymous):

thats exacly like the question says

OpenStudy (anonymous):

What does answer B say?

OpenStudy (anonymous):

-1 and 4

OpenStudy (anonymous):

there are real roots but none of A - D makes sense because i got x = -4, -1

OpenStudy (anonymous):

ok we can go to next ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

which of the following is equivalent to the equations 11-3x^2+7x-4=8x+1 A.3x^2+15x-6=0 B.3x^2+15x+16=0 C.3x^2+x-6=0 D.-3x^2+x+6=0